OpenStudy (anonymous):

Solve. (x - 5)^2 = 4. x = 3 and x = 11 x = 9 and x = 11 x = 9 and x = 1 x = 7 and x = 3

OpenStudy (anonymous):

x = 7 and x = 3

OpenStudy (mertsj):

Take the square root of both sides and get x-5 = positive or negative 2

OpenStudy (mertsj):

Then solve those two equations x-5=2 x=7 x-5=-2 x=3